### v1

# On the Robust Approximation of ASR Metrics

Abstract
Recent advances in speech foundation models are largely driven by scaling both model size and data, enabling them to perform a wide range of tasks, including speech recognition. Traditionally, ASR models are evaluated using metrics like Word Error Rate (WER) and Character Error Rate (CER), which depend on ground truth labels. As a result of limited labeled data from diverse domains and testing conditions, the true generalization capabilities of these models beyond standard benchmarks remain unclear. Moreover, labeling data is both costly and time-consuming. To address this, we propose a novel label-free approach for approximating ASR performance metrics, eliminating the need for ground truth labels. Our method utilizes multimodal embeddings in a unified space for speech and transcription representations, combined with a high-quality proxy model to compute proxy metrics. These features are used to train a regression model to predict key ASR metrics like Word Error Rate (WER) and Character Error Rate (CER). We experiment with over 40 models across 14 datasets representing both standard and in-the-wild testing conditions. Our results show that we approximate the metrics within a single-digit absolute difference across all experimental configurations, outperforming the most recent baseline by more than 50%.
On the Robust Approximation of ASR Metrics
Abdul Waheed1       Hanin Atwany2       Rita Singh1       Bhiksha Raj1 1Carnegie Mellon University      2MBZUAI {abdulw,bhiksha,rsingh}@cs.cmu.edu        hanin.atwany@mbzuai.ac.ae
1 Introduction
Automatic Speech Recognition (ASR) models have made significant advancements in recent years, achieving near-human performance on several standard evaluation benchmarks (Radford et al., 2022; Seamless Communication et al., 2023; Communication et al., 2023; Harper et al., , inter alia). These models are typically evaluated using metrics like Word Error Rate (WER) and Character Error Rate (CER) Likhomanenko et al. (2020), which are essential for assessing model performance.
However, these metrics are dependent on ground truths, which are often scarce in resource-constrained environments, and human labeling is both costly and time-consuming. To mitigate this challenge, several reference-free evaluation methods are proposed Yuksel et al. (2023b); Kalgaonkar et al. (2015); Swarup et al. (2019); Qiu et al. (2021); Del-Agua et al. (2018); Raj et al. (2011). While these approaches eliminate the reliance on labeled data, they primarily offer relative assessments of transcription quality, rather than providing precise error counts or rates. As a result, their applicability in real-world scenarios, where actionable performance metrics are crucial for further model refinement and deployment, is limited.
Given the limitations of both methods, approximating ASR metrics has emerged as a promising alternative for label-free evaluation Chowdhury and Ali (2023); Sheshadri et al. (2021b); Ali and Renals (2018). This approach typically involves training regression Jalalvand et al. (2016) and/or classification models Sheshadri et al. (2021a) on top of speech and text encoders. While this method offers a close approximation of error metrics, several important questions remain unresolved. Specifically, an approximation model trained on dataset sampled from $D$ to predict ASR metrics for a source model $M$ must be evaluated under diverse conditions: 1) on test data that is IID (independent and identically distributed) sampled from $D$; 2) on out-of-distribution (OOD) data representing diverse domains and recording conditions; 3) on IID data, but transcription from a target model $T$; and 4) on OOD data with transcriptions from a target model $T$. Most prior works Chowdhury and Ali (2023); Sheshadri et al. (2021b) focus primarily on the first condition. Moreover, recent advancements in multimodal foundation models offer new opportunities to directly train regression models on unified speech and text embeddings.
To address these critical research gaps, we propose a novel framework for approximating the performance of a wide range of ASR models, both on standard benchmarks and in-the-wild scenarios. Specifically, we compute the similarity between speech and text embeddings in a unified space, capturing the semantic alignment between the two modalities. Additionally, we incorporate a high-quality reference model as a proxy, based on the intuition that agreement with a reliable proxy correlates with transcription quality, as shown in prior works Waheed et al. (2025). These features are then used to train a regression model to predict key ASR metrics, such as WER, CER, and absolute word and character error counts.
In summary, our work represents one of the most comprehensive studies to date on approximating ASR metrics at scale, in terms of both data and model coverage. Our proposed approach serves as a reference-free evaluation particularly suited for label-scarce scenarios. Beyond evaluation, our method is especially valuable for tasks such as pseudo-labeling, where high-quality transcriptions are essential for downstream applications like knowledge distillation Waheed et al. (2024); Gandhi et al. (2023).
Our contributions are as follows:
  • We evaluate over 40 ASR models across 14 diverse evaluation setups, including both standard benchmarks and domain-specific, unseen conditions followed by training regression models to approximate ASR metrics.
  • We compare our approach with the most recent work on approximating ASR metrics and demonstrate over a 100% improvement against the strong baseline.
  • We conduct a rigorous ablation study to analyze the impact of different experimental configurations, providing deeper insights into the robustness of our approach. Our findings show that our method is resilient to diverse evaluation setups and requires only a small amount of training data.
Outline. The remainder of this paper is organized as follows: Section 2 reviews related work. Section 3 presents our proposed methodology. Sections 4 and 5 detail our experimental setup, results, and ablation study, respectively. Section 6 concludes the paper and outlines future directions.
Reproducibility. We are committed to making all code, data, and logs available upon acceptance.
2 Related Work
Automatic speech recognition (ASR) has seen remarkable progress in recent years, driven by advances in deep learning and the availability of extensive training datasets Radford et al. (2022); Communication et al. (2023). Transformer Vaswani et al. (2023) based models, in particular, have significantly contributed to these developments by effectively capturing long-range dependencies and contextual nuances in speech, achieving state-of-the-art (SOTA) performance across diverse benchmarks Kheddar et al. (2024); Dhanjal and Singh (2024); Zimerman and Wolf (2023). While traditional evaluation metrics like Word Error Rate (WER) and Character Error Rate (CER) are de-facto evaluation metrics in benchmarking ASR systems Lin et al. (2021); Park et al. (2024), scenarios where ground truth transcriptions are unavailable have caught interest in reference-free ASR evaluation methods Karbasi and Kolossa (2022); Wang et al. (2024); Kuhn et al. (2024).
Reference-free ASR evaluation methods aim to estimate ASR performance without requiring ground truth transcriptions Ospanov et al. (2024). Earlier approaches rely on heuristic features or metadata such as speaker demographics, background noise, and linguistic characteristics Litman et al. (2000); Yoon et al. (2010), limiting their applicability across varied contexts. However, recent advancements focus on deep learning-based frameworks, such as convolutional neural networks (CNNs) Elloumi et al. (2018) and contrastive learning methhods Yuksel et al. (2023a), to predict ASR quality directly from encoded speech and text. For instance, methods like NoRefER Yuksel et al. (2023b) employ Siamese architectures fine-tuned on ASR hypotheses, achieving high correlation with traditional metrics and improving WER by optimizing hypothesis ensembling Park et al. (2024).
Efforts to approximate ASR metrics explore hybrid approaches that combine traditional and reference-free methods, such as leveraging word confidence scores, linguistic embeddings, or post-processing adaptations to estimate WER and CER without explicit references Ali and Renals (2020, 2018); Kuhn et al. (2024); Negri et al. (2014). However, these approaches often suffer from reliance on specific ASR models or domain characteristics, limiting their generalizability. Unlike existing methods, our work addresses these limitations by introducing a robust, model and data-agnostic framework that evaluates ASR outputs across diverse datasets and configurations, emphasizing adaptability to unseen domains and variations.
3 Methodology
We present a scalable and robust method to approximate ASR performance metrics using multimodal unified embeddings, proxy references, and regression models. The primary goal is to eliminate reliance on ground-truth labels, enabling performance evaluation in label-scarce scenarios. The pipeline consists of three components: representation similarity in a unified speech-text embedding space, agreement with a high-quality proxy reference, and a regression model trained on these features to predict ASR metrics. Our pipeline diagram is shown in Figure 1.
3.1 Similarity in Unified Representation Space
The foundation of our approach is the SONAR model Duquenne et al. (2023), a state-of-the-art multimodal (speech-text) model trained to produce unified embeddings for both speech and text inputs. Let $x_{\text{speech}}$ represent the input speech signal and $x_{\text{text}}$ denote the corresponding ASR-generated transcription. SONAR maps these inputs to a shared embedding space, generating $e_{\text{speech}}$ and $e_{\text{text}}$:
$e_{\text{speech}}=f_{\text{SONAR}}(x_{\text{speech}}),\quad e_{\text{text}}=f_{\text{SONAR}}(x_{\text{text}})$ (1)
where $f_{\text{SONAR}}$ represents the embedding model. The alignment between these embeddings is quantified using cosine similarity:
$\text{Similarity}(x_{\text{speech}},x_{\text{text}})=\frac{e_{\text{speech}}\cdot e_{\text{text}}}{\|e_{\text{speech}}\|\|e_{\text{text}}\|}$ (2)
The similarity metric serves as an indicator of transcription quality, with higher values suggest better alignment between speech and text representations.
3.2 Agreement with a Proxy Reference
To complement the similarity score, we utilize proxy references generated by a high-quality ASR model, denoted as $x_{\text{proxy}}$. The comparison between the ASR-generated transcription $x_{\text{text}}$ and the proxy reference $x_{\text{proxy}}$ is quantified using Word Error Rate (pWER) and Character Error Rate (pCER) as defined in Appendix A.1.
These metrics assess transcription quality by comparing it with a reliable proxy reference, without using ground-truth labels at any stage. Proxy references are dynamically selected by profiling 41 models across datasets and ranking them by average performance. For each target ASR model, the reference is the highest-ranking model other than the target itself. For example, if whisper-large-v3 ranks highest, the reference for whisper will be the second-best model. This ensures the proxy reference is both relevant and reliable for evaluating the target model.
Figure 1: High-level diagram for our framework. The proxy is an ASR model that takes input speech and generates a transcription. We use the output from the source model as a hypothesis, and the output from the proxy model as a reference, to calculate metrics like WER and CER (pmetrics), which we denote as pWER and pCER. We then use this, along with the the similarity between SONAR embeddings of the input speech and the hypothesis, to train the regression that gives approximated metrics (ametrics), e.g., aWER/aCER.
3.3 Regression Model for Metric Prediction
The extracted features, including similarity scores and proxy metrics, are concatenated to form the input to a regression model. Let $z=[\text{Similarity},\text{pWER}/\text{pCER}]$ represent the feature vector. The regression model $g$ estimates the ASR metrics $\hat{y}$, denoted as aWER and/or aCER:
$\hat{y}=g(z)$ (3)
The regression model is an ensemble of Random Forest, Gradient Boosting, and Histogram-based Gradient Boosting regressors. Each base model is fine-tuned via grid search for hyperparameter optimization. The ensemble is trained to minimize the mean absolute error between predicted and ground-truth metrics. Additionally, a ridge regression model with non-negativity constraints is included in the ensemble to ensure predictions remain within valid ranges. Additional details of our regression pipeline are provided in Section 4, with hyperparameter details in Appendix A.4.
3.4 Evaluation
We evaluate the regression model’s performance across four setups, including IID and OOD data and different model configurations. Specifically, we train our regression model on one ASR system (source) on one dataset and evaluate it on both IID and OOD data for the source and target models.
Let $\mathcal{D}_{M,B}$ denote the 10 benchmark datasets, and $\mathcal{D}_{M,W}$ represent the four in-the-wild datasets, as described in Section 4.1, where $M\in\{S,T\}$ refers to either the source model $S$ or the target model $T$.
The regression model is trained on data $\mathcal{D}_{S,B}^{\text{train}}\sim\mathcal{D}_{S,B}$ and evaluated on the IID test set $\mathcal{D}_{S,B}^{\text{test-IID}}\sim\mathcal{D}_{S,B}$, consisting of $80\%$ and $20\%$ of the data, respectively. Additionally, the model is evaluated on $\mathcal{D}_{T,B}^{\text{test-IID}}$, $\mathcal{D}_{S,W}$, and $\mathcal{D}_{T,W}$. Below, we detail the formulation of each evaluation setup.
Case 1: IID Evaluation (Source $S$)
The regression model is trained on $\mathcal{D}_{S,B}^{\text{train}}$ and evaluated on $\mathcal{D}_{S,B}^{\text{test-IID}}$. Let $x_{1}^{S}=f(s,o^{S})$ represent the similarity between input speech $s$ and the ASR output $o^{S}$, and $x_{2}^{S}=g(o^{S},r)$ represent the agreement with the proxy reference $r$, where $o^{S}$ is the ASR output produced by the source model $S$. The evaluation is formulated as:
$\mathcal{L}_{\text{IID}}^{S}=\mathbb{E}_{(x_{1}^{S},x_{2}^{S},y)\sim\mathcal{D}_{S,B}^{\text{test-IID}}}\big{[}\mathcal{L}(h(x_{1}^{S},x_{2}^{S}),y)\big{]}$ (4)
Case 2: IID Evaluation (Target $T$)
The regression model trained on $\mathcal{D}_{S,B}^{\text{train}}$ is evaluated on the IID test set $\mathcal{D}_{T,B}^{\text{test-IID}}$. Let $x_{1}^{T}=f(s,o^{T})$ represent the similarity between input speech $s$ and the ASR output $o^{T}$, and $x_{2}^{T}=g(o^{T},r)$ represent the agreement with the proxy reference $r$, where $o^{T}$ is the ASR output produced by the target model $T$. The evaluation is expressed as:
$\mathcal{L}_{\text{IID}}^{T}=\mathbb{E}_{(x_{1}^{T},x_{2}^{T},y)\sim\mathcal{D}_{T,B}^{\text{test-IID}}}\big{[}\mathcal{L}(h(x_{1}^{T},x_{2}^{T}),y)\big{]}$ (5)
Case 3: OOD Evaluation (Source $S$)
The regression model trained on $\mathcal{D}_{S,B}^{\text{train}}$ is evaluated on the out-of-distribution set $\mathcal{D}_{S,W}$. Let $x_{1}^{S}=f(s,o^{S})$ represent the similarity between the input speech $s$ and the ASR output $o^{S}$, and $x_{2}^{S}=g(o^{S},r)$ represent the agreement with the proxy reference $r$, where $o^{S}$ is the ASR output produced by the source model $S$. The evaluation is defined as:
$\mathcal{L}_{\text{OOD}}^{S}=\mathbb{E}_{(x_{1}^{S},x_{2}^{S},y)\sim\mathcal{D}_{S,W}}\big{[}\mathcal{L}(h(x_{1}^{S},x_{2}^{S}),y)\big{]}$ (6)
Case 4: OOD Evaluation (Target $T$)
The regression model trained on $\mathcal{D}_{S,B}^{\text{train}}$ is evaluated on the out-of-distribution set $\mathcal{D}_{T,W}$, using the ASR output produced by the target model $T$. Let $x_{1}^{T}=f(s,o^{T})$ represent the similarity between the input speech $s$ and the ASR output $o^{T}$, and $x_{2}^{T}=g(o^{T},r)$ represent the agreement with the proxy reference $r$, where $o^{T}$ is the ASR output produced by the target model $T$. The evaluation is expressed as:
$\mathcal{L}_{\text{OOD}}^{T}=\mathbb{E}_{(x_{1}^{T},x_{2}^{T},y)\sim\mathcal{D}_{T,W}}\big{[}\mathcal{L}(h(x_{1}^{T},x_{2}^{T}),y)\big{]}$ (7)
Note. For computational feasibility, the primary experiments train the regression model on 9 out of the 10 datasets in $\mathcal{D}_{S,B}^{\text{train}}$ and evaluate it on the remaining dataset, as well as on all four datasets in $\mathcal{D}_{S,B}^{\text{OOD}}$. This process is repeated for each dataset in $\mathcal{D}_{S,B}^{\text{train}}$, ensuring robust evaluation across various testing conditions. No examples from $\mathcal{D}_{M,\text{OOD}}$ are used at any stage for training the regression model.
4 Experiments
In this section, we present the experimental setup to evaluate our ASR metrics approximation tool. We describe the datasets, models, and regression pipeline used in our experiments, highlighting the diversity of ASR systems and testing conditions.
4.1 Datasets
To evaluate the robustness and generalizability of our ASR metrics approximation tool, we use datasets sourced from multiple distributions, divided into two types: Standard Benchmark and Wild Challenge datasets. Below we describe the datasets and provide additional details in Appendix A.2 Table 3.
Standard Benchmark Datasets. We include widely used datasets representing diverse domains and acoustic conditions. LibriSpeech Panayotov et al. (2015) provides 1,000 hours of English read audiobooks, covering both clean and noisy conditions. TED-LIUM Rousseau et al. (2014) consists of TED talks from 2,000 speakers. GigaSpeech Chen et al. (2021) spans audiobooks, podcasts, and YouTube, incorporating both read and spontaneous speech. SPGISpeech Technologies (2021) features 5,000 hours of earnings calls with a focus on orthographic accuracy. Common Voice Ardila et al. (2020) is a multilingual, crowdsourced corpus with diverse accents. Earnings22 Rio et al. (2022) provides 119 hours of accented, real-world earnings calls. Additional datasets include AMI (IHM) Carletta et al. (2005), with 100 hours of English meeting recordings from non-native speakers, and People’s Speech Galvez et al. (2021), emphasizing inclusivity and linguistic diversity. SLUE-VoXCeleb Shon et al. (2022) contains conversational voice snippets, capturing diverse speaking styles and emotions.
Wild Datasets. The wild set focuses on real-world variability and challenging scenarios. Primock57 Papadopoulos Korfiatis et al. (2022) includes telemedicine consultations with diverse accents, ages, and scenarios, recorded by clinicians and actors. VoxPopuli Accented Wang et al. (2021) contains multilingual speeches from European Parliament recordings, rich in named entities. ATCOsim Jan van Doorn (2023) features 10 hours of non-native English speech from air traffic control simulations with clean utterance-level transcriptions. Additionally, we include a noisy subset of LibriSpeech Panayotov et al. (2015), which reflects challenging real-world conditions.
4.2 Models
We evaluate our approximation framework for a range of state-of-the-art ASR models, put into three categories based on their architecture and functionality. Below we describe the datasets and provide additional details in Appendix A.2 and in Table 4.
Encoder-Decoder Models. We include multiple encoder-decoder families of models capable of performing ASR tasks in a zero-shot setting. More specifically, we include whisper Radford et al. (2023) and distil-whisper Gandhi et al. (2023) models which perform really well across diverse testing settings. We also include seamless  Communication et al. (2023); Seamless Communication et al. (2023); Barrault et al. (2025), SpeechT5 Ao et al. (2022) which are unified encoder-decoder framework for tasks such as ASR, speech synthesis, translation, and voice conversion. MMS Pratap et al. (2023) supports hundreds of languages and excels in resource-constrained scenarios. Moonshine (2) Jeffries et al. (2024), a lightweight and efficient model, is designed for edge deployments with strong performance.
NeMo-ASR Models. We use multiple models from the NeMo-ASR Gulati et al. (2020); Variani et al. (2020); Noroozi et al. (2024); Tang et al. (2023); Harper et al. toolkit by NVIDIA. We include models such as Canary and Parakeet, which use highly efficient speech encoders like Fast-Conformer Rekesh et al. (2023). In addition to that, we use models based on various encoders, and decoders (CTC, RNN-T, TDT, Conformer-CTC Guo et al. (2021). In our work, we evaluate 11 models from the NeMo-ASR toolkit.
Encoder-Only and Decoder-Only Models. We include self-supervised encoder-only models and their derivatives, as well as decoder-only models like SpeechLLM. Specifically, we use Wav2Vec2 Schneider et al. (2019); Baevski et al. (2020), HuBERT Hsu et al. (2021), and Data2Vec Baevski et al. (2022). Additionally, we include speech language models like SpeechLLM Rajaa and Tushar , which combines speech embeddings with language models to predict metadata such as speaker attributes, emotions, and accents, offering robust multimodal capabilities.
4.3 Experimental Setup
We evaluate all models listed in Section 4.2 on $1000$ examples sampled randomly from the $test$ split of each dataset, as described in Section 4.1. Since all models are trained at a 16 kHz sampling rate, we (re)sample the speech accordingly. For ASR, we use greedy decoding and all other parameters are default unless otherwise specified. We apply basic text post-processing  [^1] https://bit.ly/enormwhisper before computing ASR metrics. We obtain all models from Huggingface Hub  [^2] https://huggingface.co/models and implement the ASR pipeline using the Transformers Wolf et al. (2020) library.
For multimodal embeddings, we use SONAR Duquenne et al. (2023), a 1024-dimensional sentence-level multilingual embedding model. Specifically, we utilize text_sonar_basic_encoder for text encoding and speech_sonar_basic_encoder for speech encoding. These encoders provide unified representations.
The regression framework uses a stacking ensemble with base regressors and a final estimator. Hyperparameter tuning is performed with RandomizedSearchCV to minimize MAE. The model is trained on 9 benchmark datasets and evaluated on the remaining benchmark dataset and four in-the-wild datasets. This process is repeated for all 10 benchmark datasets. Additional details of the regression pipeline are provided in Section 3 and low-level details in Appendix A.4.1.
We conduct ASR experiments on a single A100/H100 GPU, while the regression model training runs on CPUs. Although ASR time and memory consumption depend on the model size, embedding extraction for $1000$ audio-text pairs takes approximately one minute on a single consumer-grade GPU without parallelization or additional efficiency measures. Appendix A.4 provides further experimental setup details.
Baselines. Recent studies directly aligned with our approach are limited. For instance, eWER Ali and Renals (2018) and eWER2 Ali and Renals (2020) estimate error rates based on the input signal, which differs from our approach. In contrast, we incorporate the model’s output transcript into the error rate approximation function. The most closely related recent works are WERBERT Sheshadri et al. (2021a) and eWER3 Chowdhury and Ali (2023), which share a similar pipeline. Both use encoders for text, speech, and other data, followed by a regression model trained in an end-to-end setting. Since eWER3 is the more recent of the two, we use it as our baseline. In eWER3, the speech encoder is wav2vec2 Baevski et al. (2020), and the text encoder is roberta-base Liu et al. (2019), with a regression model trained on top while both encoders remain frozen. Given the unavailability of public code or pretrained models for evaluation, we implement eWER3 with some modifications to ensure a fair comparison. Specifically, we extract features from both encoders and apply PCA for dimensionality reduction on each modality before training our regression pipeline. For both speech and text, we experiment with 32 and 64 PCA components (referred to as $nc$ in Table 5.3).
5 Results
We conduct experiments using two dataset categories: standard benchmarks and in-the-wild, as described in Section 4.1. For each ASR model, a leave-one-out strategy is used, training the regression model on 9 benchmark datasets and testing it on the remaining benchmark dataset and all four in-the-wild datasets to ensure comprehensive evaluation exclusively on out-of-domain data. Additionally, in-domain testing is included in ablation studies, as detailed in Section 5.3. The regression model is trained to predict absolute error counts (word and character levels), which are normalized by the reference length to compute approximate error rates ($aWER$ and $aCER$). We also train regression models to directly predict WER and CER.
5.1 Evaluation on In-the-Wild Datasets
The wild datasets provide a realistic testbed for evaluating the regression model’s ability to approximate error rates under real-world conditions. The results are presented in Table 1. High-performing models, like canary-1b, demonstrate strong agreement between predicted and actual error rates. For example, on VP_Accented, canary-1b achieves a WER of $23.2\%$ and an $aWER$ of $12.1\%$, with a minimal difference of $1.1\%$. On Primock57, a clinical consultation dataset, the model shows robustness with a WER of $16.2\%$ and an $aWER$ of $13.4\%$, highlighting its effective generalization across diverse and domain-specific contexts.
For models like data2vec-audio-large-960h our approximation is pretty close to actual error rates with difference consistently under $2\%$ on various datasets. For example, on LibriSpeech-test-noise, the model’s actual WER is $7.2\%$ while the approximated $aWER$ is $8.6\%$, showcasing its reliability in noisy conditions. Even on acoustically complex datasets like ATCOsim, where the WER is $44.0\%$ and the $aWER$ is $51.1\%$, the model exhibits a reasonable alignment between approximated and actual error rates.
In contrast, models with high actual error rates, such as mms-1b-fl102, show slightly larger deviations, particularly on datasets with challenging conditions. For instance, on ATCOsim, the WER is $93.4\%$ and the $aWER$ is $99.0\%$, resulting in a significant deviation of $5.6\%$, the highest observed across all in-the-wild datasets. Similarly, on Primock57, where the WER is $70.2\%$ and the $aWER$ is $67.8\%$, the approximation also struggles to align due to the inherently high error rates. This highlights that extreme error cases often correspond to semantically nonsensical outputs, where the distinction between high and extremely high error rates becomes less relevant.
Model LS_Noise Primock57 ATCOsim VP_Acc
w2v2-ls 8.8/10.2 32.8/35.6 43.0/49.5 20.4/26.4
can-1b 4.1/6.4 16.2/13.4 30.4/35.5 23.2/12.1
d2v-base 14.9/16.4 39.6/41.7 66.0/71.2 28.4/33.8
d2v-large 7.2/8.6 28.3/30.7 44.0/51.1 21.4/26.5
distil-l-v2 7.3/9.2 18.3/13.0 69.5/66.7 14.9/14.5
distil-l-v3 6.1/8.3 18.4/12.9 69.0/63.6 14.8/14.0
distil-s.en 9.1/10.6 19.3/14.7 74.9/69.1 14.6/14.7
sm4t-l 11.2/12.3 41.7/37.8 75.0/82.5 29.3/19.9
sm4t-m 14.9/15.6 44.1/39.7 54.6/60.4 30.5/22.5
hub-l-ls-ft 7.3/8.8 29.5/32.0 50.4/56.9 21.4/26.6
hub-xl-ls-ft 6.8/8.3 31.1/32.9 46.7/53.0 21.8/27.7
mms-1b-a 9.5/11.1 36.2/34.4 63.4/71.8 29.9/23.8
mms-1b-f102 24.0/24.9 70.2/67.8 93.4/99.0 39.4/38.2
moon-b 11.3/12.4 19.9/18.5 65.5/66.2 17.1/20.8
moon-t 15.5/17.4 29.2/29.5 62.9/68.5 22.1/26.2
par-ctc-0.6b 4.6/7.4 16.3/13.8 32.9/42.9 16.3/13.8
par-ctc-1.1b 4.5/6.9 16.6/14.1 30.9/39.9 16.4/12.4
par-rnnt-0.6b 3.8/6.9 16.3/13.2 31.6/41.8 17.3/12.6
par-rnnt-1.1b 3.5/6.1 14.6/13.3 27.3/37.6 18.1/10.4
par-tdt-1.1b 3.4/6.0 13.5/13.2 28.3/35.7 17.9/10.2
pkt-ctc-110m 6.1/8.6 16.7/13.0 39.9/42.4 19.2/12.5
sm4t-v2-l 7.2/8.4 34.6/31.7 52.4/57.6 33.8/24.5
spchllm-1.5B 15.3/16.6 42.0/41.8 121.1/125.4 57.0/59.3
spchllm-2B 13.9/15.6 39.4/40.3 60.6/64.1 39.2/44.1
stt-cfc-l 5.8/6.8 16.1/17.6 35.9/38.0 18.6/11.5
stt-cfc-s 9.7/11.2 22.2/24.6 43.7/47.7 16.4/15.6
stt-fc-cfc-l 6.8/10.0 17.6/23.9 34.9/47.6 18.9/13.3
stt-fc-td-l 6.0/8.8 17.0/20.6 34.5/46.5 21.1/15.1
w2v2-960h 17.4/18.5 44.7/47.1 68.4/74.0 29.9/36.5
w2v2-crelpos 5.9/7.4 28.5/30.3 47.2/54.0 22.4/26.7
w2v2-crope 6.6/8.1 31.7/33.4 49.8/56.9 21.9/26.3
w2v2-l-960h 11.6/12.6 37.8/40.2 66.4/72.7 26.3/33.3
w2v2-l-lv60-s 7.8/9.4 33.1/35.5 40.5/48.8 19.3/24.9
w2v2-l-rft-ls 10.0/11.5 32.2/34.6 48.9/55.7 22.0/28.6
whisper-l 6.2/8.1 18.8/13.9 65.3/66.9 18.7/15.9
whisper-l-v2 5.4/6.6 19.0/13.1 64.8/74.8 20.0/18.1
whisper-l-v3 4.6/5.9 18.7/12.0 64.7/73.9 19.2/18.1
whisper-l-v3-t 4.9/6.0 18.5/12.3 66.0/72.5 24.3/23.2
whisper-m.en 6.5/7.9 19.5/14.0 66.2/73.8 27.6/26.4
whisper-s.en 8.2/9.7 20.0/15.1 67.1/73.8 17.3/17.5
whisper-tiny 18.5/20.7 30.0/26.6 97.6/102.5 29.8/33.2
Table 1: Actual and approximated WER ($\downarrow$), separated by a slash, on out-of-distribution wild datasets. The regression model is trained independently for each ASR model on standard benchmarks, making the wild datasets out-of-distribution. See Table 7 for full names.
5.2 Evaluation on Benchmark Datasets
We summarize results on 10 standard benchmark datasets in Appendix A.5 Tables 8 and 9. Each table reports actual WER/CER alongside the approximated WER/CER (denoted by aWER/aCER).
Figure 2: Actural and approximated WER for four models across standard benchmark.
Overall, models such as parakeet-tdt-1.1b and whisper-large-v3 show relatively small differences between WER and aWER, indicating reliable approximations. For instance, the actual WER for whisper-large-v3 on AMI_IHM is 19.0% compared to aWER 17.1%, 1.9% gap. Conversely, some challenging datasets (e.g., CV11 and Earnings22) reveal larger discrepancies for specific models, particularly those with higher overall error rates. For example, mms-1b-fl102 exhibits a wide WER/aWER gap on Earnings22, suggesting difficulty handling accented or domain-specific speech.
In general, high-performing ASR models demonstrate small WER–aWER gaps, indicating that it’s easy to approximate when error rates are low. However, models with higher WERs or faced with more acoustically or linguistically challenging test sets tend to show wider divergences. Despite these variations, most results remain within a reasonable margin, highlighting the robustness of our approximation model on diverse out-of-distribution data.
These results underscore the critical role of model quality in achieving reliable approximations. The approximation framework remains effective for high-performing models, while deviations tend to increase in cases of semantically divergent or poorly structured outputs, reflecting the inherent challenges in approximating errors for low-performing systems.
5.3 Ablation
We conduct ablation experiments to evaluate the robustness of the approximation model and the contributions of its individual components. Using the evaluation setup outlined in Section 3.4, we select data2vec-audio-base-960h as the source model ($S$) and wav2vec2-base-960h as the target model ($T$). The results are summarized in Table 5.3, where IID results correspond to Case-I 3.4, and $D$, $M$, and $D+M$ under OOD represent Case II 3.4, Case-III 3.4, and Case-IV 3.4, respectively. The reference model’s $r$ value represents the average WER across all datasets. We include reference models with varying $r$ values, such as whisper-large-v3 ($r=17.8$), whisper-medium.en ($r=20.1$), whisper-tiny ($r=33.4$), and mms-1b-fl102 ($r=51.0$).
The results in Table 5.3 demonstrate the importance of proxy references in improving the regression model’s performance. Training without proxy references (w/o PR) significantly increases the mean absolute error (MAE) across all conditions. For instance, the IID MAE increases from $1.03$ (Base) to $3.13$, and the OOD $D+M$ MAE rises from $1.07$ (Base) to $3.33$, highlighting the essential role of proxy references in approximation.
Increasing the number of high-quality proxy references (MPR) further reduces errors. Under IID conditions, the MAE decreases from $1.00$ with $n=2$ to $0.93$ with $n=5$. Similarly, in OOD $D+M$, the error drops from $1.06$ (MPR, $n=2$) to $0.95$ (MPR, $n=5$), demonstrating that multiple high-quality references enhance model robustness.
Method IID OOD
D M D + M
eWER3(nc=32) 2.030.07 2.090.04 2.06 0.03 2.120.04
eWER3(nc=64) 1.980.06 2.070.05 2.000.04 2.090.05
Base 1.030.03 1.050.01 1.030.02 1.070.01
[…]
w/o PR 3.130.07 3.220.02 3.230.05 3.330.02
[…]
w/ MPR (n=3) 0.960.02 0.970.01 0.950.02 0.990.01
w/ MPR (n=4) 0.950.02 0.960.02 0.940.02 0.980.02
w/ MPR (n=5) 0.930.02 0.930.01 0.920.02 0.950.01
w/MPR (n=10) 0.900.02 0.930.01 0.880.02 0.950.01
w/MPR (n=20) 0.890.02 0.960.02 0.870.02 0.960.02
[…]
w/ mMPR (n=5) 0.940.02 0.940.02 0.930.01 0.960.02
w/mMPR (n=10) 0.920.02 0.940.02 0.910.02 0.960.02
w/mMPR (n=20) 1.040.02 1.050.01 1.020.02 1.040.01
[…]
Base (r=20.1) 1.360.04 1.360.01 1.340.03 1.340.01
Base (r=33.4) 1.550.04 1.690.02 1.550.04 1.630.02
Base (r=51.0) 2.030.02 2.100.01 2.080.05 2.090.01
[…]
w/o S (r=20.1) 1.550.02 1.500.01 1.550.03 1.500.01
w/o S (r=33.4) 1.790.07 1.890.02 1.780.06 1.820.02
w/o S (r=51.0) 2.230.02 2.240.01 2.280.04 2.210.01
Table 2: Mean absolute error ($\downarrow$) between predicted word error count and actual error count (in absolute terms) across different configurations. PR - Proxy Reference, S - Similarity, MPR - Multiple PR, D - Data, M - Model. The OOD results are averaged across four wild datasets. n is the number of proxy references. The r ($\downarrow$) value represents the average WER for proxy reference across 14 datasets. Superscript represents the standard deviation across five runs.
The quality of references, quantified by the $r$-value, also plays a critical role. For example, in IID conditions, the MAE increases from $1.31$ for $r=17.8$ to $2.03$ for $r=51.0$. A similar trend is observed in OOD $D+M$, where the MAE rises from $1.40$ ($r=17.8$) to $2.09$ ($r=51.0$). The absence of similarity (w/o S) combined with low-quality proxies further degrades performance, underscoring the importance of both high-quality references and similarity measures. These trends are similarly observed for character-level error count approximation, as detailed in Appendix Table 6.
Figure 3: Mean absolute error ($\downarrow$) between predicted and actual word error counts across varying training data sizes for the regression model. The model was trained on 10 standard benchmarks and evaluated on four in-the-wild test sets.
Scaling Training Data for Regression.
To evaluate the impact of training data size on the regression model, we scale the data from 1K to 10K examples in increments of 1K. As shown in Figure 3, the model’s performance does not exhibit a clear trend with increasing training data size. Some datasets show slight improvements with more data; others show minimal improvement. This suggests that the regression model is largely agnostic to the size of the training data. In fact, it appears that a relatively small dataset of just 1,000 examples is sufficient to train a robust approximation model. This underscores the model’s ability to generalize effectively with limited data, making it an efficient choice for scenarios with constrained datasets.
6 Conclusion
We present a framework for approximating ASR metrics, demonstrating its effectiveness in generalizing to unseen, in-the-wild, and challenging conditions. Our results show that the model performs well with absolute error counts, consistently outperforming strong baseline, with error rates remaining relatively low. We show that our proposed method achieves consistent performance across 40 ASR models and 14 evaluation setups, including both standard benchmarks and domain-specific conditions. The trained regression model can be efficiently used to approximate ASR metrics, particularly in data-constrained environments, such as critical domains with limited labeled data. In summary, our work bridges the gap between theoretical advancements and real-world applications, paving the way for more reliable and scalable ASR systems. While in this work, we explore the impact of training data size within a single language, future work will focus on extending this framework to support multiple languages and exploring language-agnostic ASR metric approximation.
7 Limitations
In this work, we introduced a framework for approximating ASR metrics, evaluated across various ASR models and datasets. Despite the promising results, there are several limitations to consider.
Evaluation. While our evaluation setup is comprehensive, consisting of over 40 models and 14 datasets representing various acoustic and linguistic conditions such as natural noise, dialects, and accents—far surpassing previous works—we have not explored more nuanced conditions such as gender, non-native speech, and approximation across various age groups. Additionally, while the framework has demonstrated strong performance in approximating ASR metrics across multiple datasets, its generalization to highly diverse or extreme real-world conditions might still require further investigation.
Language. Furthermore, the evaluation is currently limited to a single language; expanding this framework to multiple languages or achieving language-agnostic ASR metric approximation remains an important direction for future work.
Compute. While, unlike previous works, our final approximator is a simple machine learning model that does not require GPUs to run, we do utilize a single GPU for multimodal embedding extraction, which could be performed on any consumer-grade GPU.
8 Ethics Statement
Data Collection and Release.
The datasets used in our experiments consist of publicly available ASR data from both benchmark and in-the-wild sources, as detailed in Section 4.1. We ensure that the use of these datasets aligns with the principles of fair use, specifically in a non-commercial academic context or as specified in their original license. All datasets are openly accessible, and no private or confidential data is included in this work to the best of our knowledge.
Intended Use.
By enabling the approximation of ASR performance metrics with minimal data, our work has the potential to impact applications in domains with limited data availability, such as healthcare, emergency response, and low-resource language research. We believe our approach will foster further research in scalable, low-cost ASR systems with comprehensive evaluation, benefiting industries and research areas that serve underrepresented or resource-limited populations.
Potential Misuse and Bias.
While our regression model has demonstrated effectiveness in approximating ASR metrics, it is important to consider potential misuse and bias. Given that our model is trained on diverse datasets, including those with various linguistic, acoustic, and demographic variations, there is a risk that the model may inherit biases present in the data, particularly with respect to accents, dialects, and socio-linguistic factors. Additionally, as our model approximates error rates, it could be misused in applications where the approximation may not be sufficient for real-world critical tasks. We recommend cautious deployment and further evaluation in sensitive applications, especially those where fairness and accuracy are critical.
Appendix A Appendix
A.1 Methodology
$\text{pWER}(x_{\text{text}},x_{\text{proxy}})$ $=\frac{\text{EditDistance}(x_{\text{text}},x_{\text{proxy}})}{\text{WordCount}(x_{\text{proxy}})}$ (8)
$\text{pCER}(x_{\text{text}},x_{\text{proxy}})$ $=\frac{\text{EditDistance}(x_{\text{text}},x_{\text{proxy}})}{\text{CharCount}(x_{\text{proxy}})}$ (9)
A.2 Datasets
To evaluate the robustness and generalizability of our ASR metrics approximation tool, data were sourced from multiple repositories, which we divided into two distinct groups: Standard Benchmark and Wild Challenge dataset.
A.2.1 Standard Benchmark Datasets
There are six datasets in total that fall under the benchmark group. These datasets are categorized based on their frequent use in ASR model training and their representation of commonly encountered domains in real-world applications.
LibriSpeech Panayotov et al. (2015). prioritized speaker and content balance over explicit consideration of speech characteristics. It comprises approximately 1000 hours of English read audiobooks, with subsets featuring both clean and noisy speech conditions to simulate different acoustic environments. While the dataset covers diverse subject matter, its focus on formal, clear speech from public domain books means it lacks the natural variability of spontaneous speech, limiting its representation of conversational or informal dialogue. TED-LIUM Rousseau et al. (2014). contains TED Talks totaling 452 hours of English speech data from approximately 2,000 speakers, recorded in close-talk microphone conditions. The corpus features narrated speaking styles, capturing clear and articulate speech. While it provides non-orthographic transcriptions, lacking formatting such as capitalization and punctuation, it remains a valuable resource for training and benchmarking automatic speech recognition (ASR) models. GigaSpeech Chen et al. (2021). is a multi-domain, multi-style speech recognition corpus incorporating diverse acoustic and linguistic conditions. It sources audio from three primary domains: audiobooks, podcasts, and YouTube, covering a wide range of speaking styles, including both read and spontaneous speech. The dataset covers a broad spectrum of topics, such as arts, science, sports, and more, making it highly versatile for training robust speech recognition models. SPGISpeech Technologies (2021). contains 5,000 hours of professionally transcribed audio from corporate earnings calls, featuring both spontaneous and narrated speaking styles. It emphasizes orthographic accuracy, providing fully formatted text with capitalization, punctuation, and denormalization of non-standard words. Common Voice Ardila et al. (2020). (a multilingual corpus of narrated prompts built through crowdsourcing. Recorded in teleconference conditions, the corpus features narrated speaking styles and emphasizes inclusivity by covering a wide range of accents and languages, including low-resource ones. Earnings22 Rio et al. (2022). is a 119-hour corpus of English-language earnings calls from global companies, designed to address the lack of real-world, accented speech data in ASR benchmarking
AMI (IHM) Carletta et al. (2005). The AMI Meeting Corpus is a 100-hour dataset of English meeting recordings, featuring multimodal data synchronized across close-talking and far-field microphones, room-view and individual cameras, slide projectors, and whiteboards. It includes mostly non-native speakers recorded in three rooms with varying acoustics. Digital pens capture unsynchronized handwritten notes, supporting research in speech recognition, diarization, and multimodal interaction. Available under edinburghcstr/ami, it is widely used for meeting analysis and speech processing studies.
People’s Speech Galvez et al. (2021). Thousands of hours of labeled speech data collected from diverse speakers, covering a wide range of topics, accents, and speaking styles. The dataset emphasizes inclusivity and linguistic diversity, making it suitable for developing robust and generalized speech models. It is widely used in academic and industrial research to advance the state-of-the-art in automatic speech recognition (ASR) and other speech-related applications.
SLUE - VolxCeleb Shon et al. (2022).consists of single-sided conversational voice snippets extracted from YouTube videos, originally designed for speaker recognition. The dataset represents natural, unscripted speech in diverse real-world settings, capturing a wide range of speaking styles, emotions, and acoustic conditions. Utterances containing slurs were excluded, and partial words were trimmed using a forced aligner to ensure clean, usable segments.
A.2.2 Wild Challenge Set
Primock57 Papadopoulos Korfiatis et al. (2022). contains mock consultations conducted by seven clinicians and 57 actors posing as patients, representing a diverse range of ethnicities, accents, and ages. Each actor was provided with a detailed case card outlining a primary care scenario, such as urinary tract infections, cardiovascular issues, or mental health concerns, ensuring the conversations were realistic and clinically relevant. The consultations were recorded using telemedicine software, capturing separate audio channels for clinicians and patients, and transcribed by experienced professionals to ensure accuracy. VoxPopuli Accented Wang et al. (2021). is a comprehensive multilingual speech corpus derived from European Parliament event recordings. It includes audio, transcripts, and timestamps sourced directly from the official Parliament website. Due to its origin, the dataset features a rich collection of named entities, making it particularly suitable for tasks like Named Entity Recognition (NER).
ATCOsim Jan van Doorn (2023).is a specialized database containing ten hours of English speech from ten non-native speakers, recorded during real-time ATC simulations using close-talk headset microphones. It features orthographic transcriptions, speaker metadata, and session details. With a 32 kHz sampling frequency and 10,078 clean, utterance-level recordings.
Name Type Description
LibriSpeech Bench A corpus of approximately 1,000 hours of 16kHz read English speech, derived from LibriVox audiobooks, segmented and aligned for ASR tasks.
TED-LIUM Bench Contains TED Talks totaling 452 hours of English speech data from approximately 2,000 speakers, recorded in close-talk microphone conditions.
GigaSpeech Bench A multi-domain, multi-style speech recognition corpus incorporating diverse acoustic and linguistic conditions, sourced from audiobooks, podcasts, and YouTube.
SPGISpeech Bench Contains 5,000 hours of professionally transcribed audio from corporate earnings calls, featuring both spontaneous and narrated speaking styles.
Common Voice Bench A multilingual corpus of narrated prompts built through crowdsourcing, recorded in teleconference conditions, covering a wide range of accents and languages.
Earnings22 Bench A 119-hour corpus of English-language earnings calls from global companies, designed to address the lack of real-world, accented speech data in ASR benchmarking.
AMI (IHM) Bench The AMI Meeting Corpus is a 100-hour dataset of English meeting recordings, featuring multimodal data synchronized across various devices.
People’s Speech Bench Contains thousands of hours of labeled speech data collected from diverse speakers, covering a wide range of topics, accents, and speaking styles.
SLUE - VoxCeleb Wild Consists of single-sided conversational voice snippets extracted from YouTube videos, originally designed for speaker recognition.
Primock57 Wild Contains mock consultations conducted by seven clinicians and 57 actors posing as patients, representing a diverse range of ethnicities, accents, and ages.
VoxPopuli Accented Wild A comprehensive multilingual speech corpus derived from European Parliament event recordings, featuring a rich collection of named entities.
ATCOsim Wild A specialized database containing ten hours of English speech from ten non-native speakers, recorded during real-time air traffic control simulations.
Table 3: Overview of various ASR along with brief description.
A.3 Models
Whisper Models Radford et al. (2023). is a transformer-based model that processes 80-dimensional log-mel filter bank features from 16 kHz audio, utilizing a 2D CNN stack followed by a transformer encoder-decoder architecture. Trained on a vast multilingual dataset of 680,000 hours, it incorporates timestamp tokens into its vocabulary and operates on 30-second audio windows during inference, auto-regressively generating text sequences while leveraging encoder outputs as context. Variants of Whisper, such as Distilled, Large, Base, and Medium, offer different trade-offs in model size and performance, catering to diverse computational and accuracy requirements.
Seamless Models Communication et al. (2023); Seamless Communication et al. (2023); Barrault et al. (2025). is a cutting-edge multilingual and multitask model for speech and text translation. Built on the UnitY architecture, it uses w2v-BERT 2.0 for speech encoding and NLLB for text encoding, supporting nearly 100 languages. A text decoder handles ASR and translation, while a text-to-unit (T2U) model and multilingual HiFi-GAN vocoder generate speech. Leveraging SONAR embeddings and SeamlessAlign (443,000 hours of aligned speech/text data), it achieves SOTA results in ASR, speech-to-text, speech-to-speech, and text-to-text translation, excelling in low-resource languages. It introduces BLASER 2.0 for robust evaluation and outperforms competitors in noisy environments.
Nemo-ASR-Models Gulati et al. (2020); Variani et al. (2020); Rekesh et al. (2023); Noroozi et al. (2024); Tang et al. (2023); Harper et al. We included several NVIDIA’s NeMo advanced automatic speech recognition (ASR) models, including Canary, Parakeet (110M, 0.6B, and 1.1b), Conformer-CTC, and Fast-Conformer, as each is designed for specific use cases and optimized for performance. Canary-1B is a state-of-the-art multilingual, multitask model featuring a FastConformer encoder and Transformer decoder. The Parakeet family includes models with a FastConformer encoder paired with different decoders: CTC, RNN-T, or TDT. Conformer-CTC is a non-autoregressive model based on the Conformer architecture, combining self-attention and convolution for global and local feature extraction. It uses CTC loss and a linear decoder, supporting both sub-word (BPE) and character-level encodings. While Fast-Conformer is an optimized version of the Conformer architecture, offering significant speed improvements (2.4x faster) with minimal quality degradation. It uses 8x depthwise convolutional subsampling and reduced kernel sizes for efficiency.
Wav2Vec2 Models Schneider et al. (2019); Baevski et al. (2020). is a self-supervised pre-trained model designed to process raw audio inputs and generate speech representations. The model architecture consists of three key components: a convolutional feature encoder, a context network, and a quantization module. The convolutional feature encoder converts raw waveforms into latent representations, which are then processed by the context network a transformer based stack with 24 blocks, a hidden size of 1024, 16 attention heads, and a feed-forward dimension of 4096 to capture contextual information.The quantization module maps these latent representations to quantized forms.
HuBERT Models Hsu et al. (2021). is a self-supervised learning framework designed for speech representation learning where CNN-encoded audio features are randomly masked. During training, the model predicts cluster assignments for masked regions of the input speech, forcing it to learn both acoustic and language models from continuous inputs.
Audio/Speech Language Models 1.5B and 2B  Rajaa and Tushar is a multi-modal Language Model designed to analyze and predict metadata from a speaker’s turn in a conversation. It integrates a speech encoder to convert speech signals into meaningful embeddings, which are then processed alongside text instructions by TinyLlama-1.1B-Chat-v1.0 to generate predictions. The model accepts 16 KHz audio inputs and predicts metadata such as SpeechActivity, Transcript, Gender, Age, Accent, and Emotion.
SpeechT5 Ao et al. (2022). unified modal framework capable of handling a wide range of tasks, including automatic speech recognition (ASR), speech synthesis, speech translation, voice conversion, speech enhancement, and speaker identification.Its audio post-net, which can incorporate speaker embeddings to enable prosody transfer, making it effective for tasks like voice conversion and speech synthesis. By leveraging its encoder-decoder architecture, SpeechT5 can generate high-quality mel-spectrograms from text input while preserving speaker-specific characteristics like emotion and gender.
Model Type and Models Description
nemo_asr
– parakeet-ctc-1.1b
– parakeet-ctc-0.6b
– stt_en_conformer_ctc_large
– stt_en_fastconformer_ctc_large
– stt_en_conformer_ctc_small
– parakeet-tdt-1.1b
– parakeet-rnnt-1.1b
– parakeet-rnnt-0.6b
– stt_en_fastconformer_transducer_large
– parakeet-tdt_ctc-110m
– canary-1b
NVIDIA’s NeMo ASR models offer diverse architectures for speech-to-text applications. The Conformer-CTC model combines self-attention and convolutional operations, using Connectionist Temporal Classification (CTC) loss for efficient transcription. The Conformer-Transducer extends this by incorporating a Recurrent Neural Network Transducer (RNNT) decoder for autoregressive modeling. The Conformer-HAT variant separates label and blank score predictions, enhancing integration with external language models. For improved performance, the Fast-Conformer introduces depthwise convolutional subsampling, achieving approximately 2.4x faster encoding with minimal accuracy loss.
speechbrain
– asr-wav2vec2-librispeech
SpeechBrain provides robust models for ASR and speaker recognition.
data2vec
– data2vec-audio-large-960h
– data2vec-audio-base-960h
Data2Vec models by Facebook are designed for speech representation learning and ASR. These models use a unified learning framework for multiple modalities.
wav2vec2
– wav2vec2-large-960h-lv60-self
– wav2vec2-large-robust-ft-libri-960h
– wav2vec2-large-960h
– wav2vec2-base-960h
– wav2vec2-conformer-rope-large-960h-ft
– wav2vec2-conformer-rel-pos-large-960h-ft
Wav2Vec2 models leverage self-supervised learning on raw audio for ASR. With advanced configurations, these models provide high accuracy for diverse speech-to-text tasks.
mms
– mms-1b-all
– mms-1b-fl102
The Multilingual Speech (MMS) models by Facebook excel at speech recognition for multiple languages and accents.
hubert
– hubert-xlarge-ls960-ft
– hubert-large-ls960-ft
HuBERT models provide high-quality speech representations for ASR and other downstream speech tasks.
seamless
– hf-seamless-m4t-large
– hf-seamless-m4t-medium
– seamless-m4t-v2-large
Seamless models focus on multilingual transcription and translation, offering robust real-time speech processing solutions.
speechllm
– speechllm-1.5B
– speechllm-2B
SpeechLLM models are fine-tuned for ASR and text generation, leveraging billions of parameters for high performance.
whisper
– whisper-large-v3
– distil-large-v3
– whisper-large-v2
– whisper-large-v3-turbo
– distil-large-v2
– whisper-large
– whisper-tiny
– whisper-medium.en
– distil-small.en
– whisper-small.en
Whisper models by OpenAI provide state-of-the-art transcription and translation capabilities for multilingual ASR. These models range from tiny to large configurations.
moonshine
– moonshine-base
– moonshine-tiny
Moonshine models are lightweight and optimized for efficient ASR on edge devices with minimal computational resources.
Table 4: Overview of various ASR along with brief description.
A.4 Experiments
A.4.1 Regression Pipeline.
The regression framework is a stacking ensemble comprising multiple base regressors and a final estimator. We perform basic hyperparameter tuning using RandomizedSearchCV with 5-fold cross-validation, with the objective to minimize mean absolute error (MAE). The search explores key hyperparameters such as n_estimators, max_depth, learning_rate, and min_samples_split, balancing model complexity and generalization. We provide hyperparameter and other details in 5. The model is trained on 14 datasets divided into two groups: bench (10 standard benchmark datasets) and in-the-wild (4 diverse, real-world datasets). A leave-one-out strategy is applied to the bench set, where the model is trained on 9 datasets and evaluated on the remaining one. All trained models are also evaluated on the in-the-wild set, which remains isolated during training to assess out-of-domain generalization.
Model Hyperparameter Values
Random Forest (RF) n_estimators {100, 200, 300, 500, 700, 1000}
max_depth {5, 10, 15, 20, 25, 30}
min_samples_split {2, 5, 10, 15, 20}
min_samples_leaf {1, 2, 4, 8}
Gradient Boosting (GBR) n_estimators {100, 200, 400, 600, 800}
learning_rate {0.001, 0.01, 0.05, 0.1, 0.2}
max_depth {3, 5, 7, 10}
min_impurity_decrease {0.0, 0.001, 0.01, 0.1, 0.2}
HistGradientBoosting (HGB) max_iter {100, 200, 300, 400, 500}
learning_rate {0.001, 0.01, 0.05, 0.1, 0.2}
max_depth {3, 5, 7, 10, 15}
loss {Poisson}
Ridge Regression (Final Estimator) alpha {1e-3, 1e-2, 0.1, 1, 10, 100, 1000}
positive {True}
Pipeline passthrough {True}
Table 5: Hyperparameter details for regression model.
A.5 Results
Method IID OOD
D M D + M
Base 3.790.16 3.560.06 3.760.18 3.690.06
w/o S 3.830.14 3.650.06 3.820.16 3.730.07
w/o PR 8.430.28 8.360.08 8.670.24 8.660.08
w/ MPR (n=2) 3.690.14 3.570.06 3.660.17 3.690.06
w/ MPR (n=3) 3.620.13 3.440.07 3.580.15 3.560.07
w/ MPR (n=4) 3.570.13 3.400.06 3.530.13 3.520.06
w/ MPR (n=5) 3.490.13 3.370.06 3.470.12 3.490.07
w/ mMPR (n=3) 3.610.15 3.400.09 3.570.13 3.510.09
w/ mMPR (n=5) 3.800.15 3.470.03 3.770.13 3.560.04
Base (r=11.9) 4.680.17 5.160.06 4.640.16 5.060.05
Base (r=14.0) 4.840.18 4.880.07 4.750.17 4.770.07
Base (r=20.2) 5.130.12 5.380.07 5.120.10 5.300.07
Base (r=23.5) 5.600.13 6.120.07 5.690.21 6.030.05
w/o S (r=11.9) 5.500.21 5.840.06 5.550.21 5.650.05
w/o S (r=14.0) 5.730.12 5.500.05 5.710.13 5.370.06
w/o S (r=20.2) 6.160.18 6.240.08 6.130.10 5.970.09
w/o S (r=23.5) 6.380.09 6.770.08 6.430.16 6.580.08
Table 6: Mean absolute error between predicted character error count and actual character error count (in absolute terms) across different configurations. R - Regression, C - Classification, PR - Proxy Reference, S - Silarity, MPR - Multiple PR. The OOD results are averaged across five wild datasets. Superscript represents the standard deviation across five runs.
Model LS_Noise Primock57 Atcosim VP_accented
asr-wav2vec2-librispeech 4.2/5.8 17.2/20.8 18.8/21.9 9.8/14.0
canary-1b 1.5/3.8 10.1/9.7 16.4/19.4 15.6/9.0
data2vec-audio-base-960h 7.0/8.1 20.5/23.7 29.5/32.0 13.3/17.8
data2vec-audio-large-960h 3.1/4.2 14.1/17.4 20.0/23.8 10.6/14.4
distil-large-v2 3.5/5.2 11.5/9.2 49.5/41.8 10.2/9.4
distil-large-v3 2.7/4.6 11.9/9.1 49.4/40.5 10.1/9.0
distil-small.en 4.2/5.8 12.2/10.4 50.7/41.8 9.7/9.2
hf-seamless-m4t-large 6.5/7.5 32.1/30.6 54.7/57.2 21.8/15.8
hf-seamless-m4t-medium 9.4/10.1 34.4/32.7 35.5/37.9 23.1/17.9
hubert-large-ls960-ft 3.0/4.2 14.4/17.4 21.3/25.0 10.0/14.3
hubert-xlarge-ls960-ft 2.7/4.1 15.3/18.1 20.1/23.8 10.2/14.5
mms-1b-all 3.6/4.8 19.5/19.1 27.2/31.8 17.0/12.6
mms-1b-fl102 9.0/10.0 35.0/33.2 55.4/57.3 18.2/17.6
moonshine-base 5.7/6.8 12.4/12.1 42.6/39.5 10.9/12.6
moonshine-tiny 8.5/9.9 17.9/19.0 38.2/38.4 13.2/15.1
parakeet-ctc-0.6b 1.7/3.7 10.1/9.9 16.2/22.7 9.7/9.0
parakeet-ctc-1.1b 1.7/3.6 10.0/10.1 14.8/21.4 10.0/8.0
parakeet-rnnt-0.6b 1.3/3.4 10.1/9.4 16.9/24.1 10.9/8.8
parakeet-rnnt-1.1b 1.3/3.3 9.1/9.7 14.5/21.3 11.2/7.2
parakeet-tdt-1.1b 1.1/3.1 8.2/9.4 14.0/20.0 10.9/6.8
parakeet-tdt_ctc-110m 2.5/4.7 10.3/9.2 22.3/24.2 12.4/8.6
seamless-m4t-v2-large 3.5/4.6 24.6/23.7 31.6/35.8 25.2/19.8
speechllm-1.5B 9.9/11.2 30.1/31.4 85.4/88.7 47.3/49.0
speechllm-2B 8.4/9.3 25.3/27.7 33.5/36.1 24.0/28.3
stt_en_conformer_ctc_large 2.1/3.4 8.8/11.2 17.1/18.2 11.1/7.5
stt_en_conformer_ctc_small 4.3/5.7 12.7/15.6 21.6/23.6 9.5/9.7
stt_en_fastconformer_ctc_large 3.0/5.6 10.1/16.3 17.3/25.1 11.5/9.2
stt_en_fastconformer_transducer_large 2.8/5.0 10.6/14.3 18.7/25.3 14.2/11.9
wav2vec2-base-960h 7.9/9.1 23.3/26.7 30.3/33.2 13.7/18.9
wav2vec2-conformer-rel-pos-large-960h-ft 2.6/3.8 14.7/17.4 21.0/24.5 11.2/14.7
wav2vec2-conformer-rope-large-960h-ft 2.9/4.0 16.1/18.7 22.2/25.9 11.0/14.2
wav2vec2-large-960h 5.1/6.3 19.1/22.4 28.8/31.8 12.2/17.4
wav2vec2-large-960h-lv60-self 3.5/5.0 17.6/21.0 18.6/23.0 9.3/13.6
wav2vec2-large-robust-ft-libri-960h 4.5/5.8 15.7/19.0 20.7/24.2 10.0/14.7
whisper-large 2.9/4.2 13.7/10.6 49.3/47.5 13.7/11.9
whisper-large-v2 2.6/3.8 15.3/12.5 48.6/51.5 15.3/14.2
whisper-large-v3 2.0/3.3 12.3/8.7 48.9/48.3 14.3/13.6
whisper-large-v3-turbo 2.0/3.2 12.4/8.8 48.3/49.9 19.7/19.0
whisper-medium.en 3.3/4.3 13.1/10.5 49.1/49.2 23.8/20.6
whisper-small.en 4.2/5.3 13.1/10.8 48.4/51.2 12.5/12.7
whisper-tiny 9.8/11.3 19.3/18.2 60.8/63.0 21.0/22.3
Table 7: Actual and approximated CER ($\downarrow$), separated by a slash, on out-of-distribution wild datasets. The regression model is trained independently for each ASR model on standard benchmarks, making the wild datasets out-of-distribution.
Figure 4: Comparison of Actual vs Approximated WER across models.
Figure 5: Comparison of Actual vs Approximated CER across models.
Model AMI_IHM CV11 Earnings22 Gigaspeech LibriSpeech_clean
WER/aWER CER/aCER WER/aWER CER/aCER WER/aWER CER/aCER WER/aWER CER/aCER WER/aWER CER/aCER
asr-wav2vec2-librispeech 28.4/30.5 13.8/17.6 25.0/29.7 11.7/15.0 37.3/33.2 21.3/16.1 16.6/16.5 6.9/7.4 1.8/3.8 0.5/2.2
canary-1b 15.4/17.6 9.2/12.7 8.7/14.2 4.1/8.5 21.8/16.0 15.8/9.1 11.1/6.9 5.5/4.3 1.5/5.7 0.5/3.5
data2vec-audio-base-960h 39.9/40.4 19.9/23.5 37.8/42.3 18.3/21.7 50.8/48.6 28.0/25.0 23.8/23.5 10.1/10.8 2.8/4.0 0.9/1.6
data2vec-audio-large-960h 34.1/36.1 16.9/21.2 23.3/27.9 10.9/14.1 37.7/34.5 21.2/16.7 17.0/16.6 7.2/7.4 1.8/3.9 0.5/1.7
distil-large-v2 17.8/16.8 11.2/11.5 14.2/19.7 7.1/10.6 19.3/20.0 12.5/13.7 12.8/8.2 7.1/5.4 3.4/6.7 1.5/4.2
distil-large-v3 18.5/17.3 11.6/11.7 13.7/19.4 6.6/10.3 18.4/19.8 12.1/13.0 12.2/7.9 6.9/5.3 2.8/6.6 1.2/4.1
distil-small.en 18.5/18.4 11.1/12.6 18.5/23.1 9.4/12.5 21.2/21.4 13.6/14.7 13.1/8.6 7.3/5.7 3.7/7.6 1.6/4.5
hf-seamless-m4t-large 36.3/33.9 25.4/25.1 9.5/13.2 5.1/7.4 30.7/32.8 21.1/23.9 24.2/21.1 16.7/15.7 3.2/4.8 1.5/2.7
hf-seamless-m4t-medium 40.6/37.2 29.5/28.9 11.3/14.3 6.0/7.4 33.7/35.9 23.9/26.4 30.2/28.1 22.3/21.7 3.8/5.3 1.6/2.9
hubert-large-ls960-ft 31.1/33.6 15.2/19.8 24.1/28.8 10.6/13.6 37.6/34.4 20.6/16.3 19.3/18.3 8.1/7.8 2.1/3.7 0.6/1.6
hubert-xlarge-ls960-ft 31.1/34.3 15.0/20.0 24.1/28.7 10.5/13.9 37.3/34.9 20.4/15.9 18.1/17.4 7.3/7.6 2.0/3.8 0.6/1.7
mms-1b-all 37.0/36.2 19.1/20.8 22.5/27.5 8.9/12.5 34.1/30.6 19.6/15.1 19.4/16.9 8.3/7.6 4.2/6.2 1.3/2.7
mms-1b-fl102 75.4/73.3 35.1/33.9 42.6/45.3 17.8/19.9 50.6/52.3 24.2/26.5 37.2/35.7 15.7/15.2 15.8/17.3 5.1/5.9
moonshine-base 15.6/24.7 9.4/16.7 20.8/25.4 10.8/13.8 24.3/25.6 15.9/16.6 14.2/10.4 8.1/6.8 3.4/6.3 1.3/3.7
moonshine-tiny 21.3/25.3 12.8/16.7 26.7/31.7 14.4/17.3 31.2/32.7 19.7/20.2 16.6/14.1 9.1/8.6 4.5/7.2 1.8/4.2
parakeet-ctc-0.6b 17.0/23.1 10.0/16.3 10.7/21.1 5.1/11.2 24.7/19.1 16.9/11.5 12.0/8.6 6.1/5.2 2.0/5.1 0.7/2.5
parakeet-ctc-1.1b 15.7/21.4 9.0/15.3 10.5/20.1 5.2/11.0 24.0/17.7 16.6/10.7 12.2/7.9 6.2/5.0 1.8/5.4 0.5/2.6
parakeet-rnnt-0.6b 18.8/24.0 11.7/17.9 8.5/19.9 4.2/10.8 25.2/18.7 17.5/11.5 11.7/9.0 6.2/5.4 1.8/5.5 0.6/3.2
parakeet-rnnt-1.1b 18.6/23.5 11.7/17.2 6.7/19.6 3.4/10.5 25.7/17.9 18.4/11.4 11.3/8.4 6.0/5.0 1.5/5.0 0.5/3.3
parakeet-tdt-1.1b 17.1/23.5 10.2/16.9 7.2/19.6 3.4/10.6 24.5/16.6 17.1/10.0 10.2/7.8 4.9/4.7 1.3/6.0 0.4/2.9
parakeet-tdt_ctc-110m 18.5/18.8 10.7/13.6 12.7/17.7 6.9/10.1 22.2/14.8 15.7/9.2 12.6/8.2 6.2/5.0 2.6/6.7 0.9/3.8
seamless-m4t-v2-large 43.0/42.3 30.2/30.2 8.2/12.3 3.9/6.3 47.3/47.4 33.7/33.9 25.7/23.2 18.1/17.2 2.7/4.4 1.0/2.5
speechllm-1.5B 67.7/69.3 51.5/55.0 18.5/22.7 10.0/12.7 50.8/48.2 38.3/35.4 27.5/26.0 18.1/18.3 10.5/12.1 7.3/9.2
speechllm-2B 38.6/40.8 24.3/28.2 24.6/28.2 16.5/18.3 47.3/45.0 32.5/30.8 24.4/23.6 13.5/13.8 7.0/9.3 4.5/4.8
stt_en_conformer_ctc_large 15.3/19.9 7.9/13.4 10.4/15.4 4.7/8.0 24.8/20.0 16.4/10.7 13.2/10.6 5.9/5.6 2.2/3.7 0.7/2.4
stt_en_conformer_ctc_small 21.2/24.4 11.2/15.4 19.1/24.1 8.9/12.2 29.3/25.3 19.0/14.1 15.5/14.9 7.2/7.7 3.9/5.4 1.4/3.1
stt_en_fastconformer_ctc_large 20.3/24.0 11.7/15.3 9.5/19.3 4.6/10.2 27.3/21.5 18.3/13.0 14.5/14.7 7.2/8.2 1.9/5.2 0.7/2.8
stt_en_fastconformer_transducer_large 19.8/22.0 12.9/16.8 9.3/18.0 4.7/9.8 31.5/26.9 23.0/18.8 13.6/13.2 7.4/7.8 1.8/3.9 0.6/2.6
wav2vec2-base-960h 37.9/38.7 18.7/21.9 40.6/45.7 19.5/22.8 51.1/48.6 28.2/25.4 26.2/26.6 11.7/12.2 3.7/4.5 1.1/1.9
wav2vec2-conformer-rel-pos-large-960h-ft 35.0/38.7 18.5/24.1 23.7/28.0 10.7/13.7 38.4/36.2 21.7/17.6 18.5/17.2 8.5/7.9 1.6/3.3 0.5/1.5
wav2vec2-conformer-rope-large-960h-ft 34.3/36.4 18.0/22.7 23.6/28.5 11.6/15.0 36.9/33.9 21.4/16.7 17.9/17.7 7.3/7.6 1.8/3.8 0.5/1.6
wav2vec2-large-960h 34.0/36.4 16.4/20.2 34.1/38.6 16.2/19.4 46.4/43.4 25.4/21.7 20.6/20.5 8.6/9.1 2.9/4.3 0.8/2.1
wav2vec2-large-960h-lv60-self 29.1/31.5 15.5/19.5 23.1/28.8 11.0/15.0 36.7/32.5 20.8/15.7 17.6/17.2 7.5/8.0 1.7/3.5 0.5/1.9
wav2vec2-large-robust-ft-libri-960h 30.5/33.9 13.8/19.2 25.0/29.3 10.7/13.8 37.1/33.5 20.5/15.7 18.0/17.6 7.1/7.7 2.8/4.3 0.8/2.3
whisper-large 18.5/18.3 12.3/13.0 13.0/18.0 6.6/9.3 18.8/20.3 12.3/14.9 12.2/7.7 7.1/5.1 2.8/5.1 1.4/3.5
whisper-large-v2 18.6/17.1 12.1/11.8 11.3/15.5 5.7/8.0 19.0/21.5 13.0/15.8 12.5/7.1 7.3/4.9 2.8/5.1 1.5/3.2
whisper-large-v3 19.0/17.1 12.3/12.0 9.9/14.5 4.9/6.9 18.2/20.7 12.1/14.9 12.5/7.3 7.2/4.9 2.2/4.0 0.9/2.9
whisper-large-v3-turbo 19.0/17.5 12.3/11.9 12.6/16.1 6.3/8.1 18.8/21.1 12.9/15.6 12.2/6.8 7.1/4.6 2.4/4.4 1.1/2.5
whisper-medium.en 20.3/18.8 13.7/14.5 14.3/17.8 7.2/9.2 20.1/22.6 13.3/16.5 12.8/7.6 7.6/5.5 3.3/5.5 1.8/3.5
whisper-small.en 19.8/17.9 12.8/12.2 17.8/21.9 9.3/11.3 20.6/22.9 13.6/16.3 12.8/8.1 7.3/5.1 3.3/5.6 1.4/3.1
whisper-tiny 26.7/25.1 16.7/17.0 33.5/40.3 17.7/20.9 33.8/35.4 22.0/25.4 20.6/18.2 12.0/11.0 7.9/11.2 3.4/5.1
Table 8: Actual and approximated WER and CER, separated by a slash, across five standard datasets. The regression model is trained on nine datasets and tested on one, with this process repeated for all datasets, ensuring that the test data is always out-of-distribution.
Model peoples_speech slue_voxceleb spgispeech_S tedlium-dev-test voxpopuli_en
WER/aWER CER/aCER WER/aWER CER/aCER WER/aWER CER/aCER WER/aWER CER/aCER WER/aWER CER/aCER
asr-wav2vec2-librispeech 35.6/32.9 19.8/17.7 19.5/20.4 9.8/12.2 11.1/12.2 4.8/4.7 10.3/11.1 5.2/5.8 14.3/12.6 6.6/5.1
canary-1b 16.5/22.5 11.1/15.2 14.9/11.1 10.8/8.2 3.2/6.7 2.0/3.9 7.9/7.6 5.9/5.0 6.4/4.9 3.9/3.4
data2vec-audio-base-960h 43.4/38.6 24.4/20.8 26.1/27.6 13.0/15.5 19.2/19.8 8.2/7.9 13.6/14.2 6.3/6.4 18.9/17.5 8.5/7.1
data2vec-audio-large-960h 35.1/31.3 20.0/17.3 20.4/22.1 10.3/12.9 11.3/12.0 4.9/4.7 9.9/10.6 4.5/5.0 14.9/13.4 6.9/5.5
distil-large-v2 17.4/21.8 12.2/14.1 16.0/10.8 11.4/7.4 3.7/7.6 1.8/4.5 10.4/8.5 8.8/5.4 9.5/8.2 5.8/4.6
distil-large-v3 17.4/21.6 12.4/13.8 14.4/10.0 10.3/6.8 3.6/7.4 1.8/4.5 10.7/9.2 8.6/5.7 9.3/6.7 5.8/4.1
distil-small.en 19.0/22.5 13.3/14.3 15.9/11.4 11.3/7.8 4.0/7.9 1.9/4.7 10.8/8.8 9.1/5.6 10.2/7.4 6.4/4.3
hf-seamless-m4t-large 38.5/41.5 29.2/30.1 47.2/42.8 39.4/36.1 16.2/18.7 11.5/13.0 19.8/19.1 15.7/14.4 8.1/6.5 5.0/3.6
hf-seamless-m4t-medium 43.6/45.7 33.6/34.2 50.9/47.4 43.2/40.3 12.9/15.5 8.8/10.4 27.0/26.2 21.3/20.0 8.8/7.3 5.5/4.4
hubert-large-ls960-ft 34.1/31.3 18.8/17.7 20.8/22.0 10.1/12.3 11.6/12.4 4.9/4.6 11.0/12.0 5.3/5.4 15.0/13.6 6.9/5.4
hubert-xlarge-ls960-ft 35.5/31.5 19.5/16.0 20.3/22.1 9.9/12.3 11.9/12.3 4.8/4.5 10.1/11.2 4.2/5.0 14.5/12.8 6.7/5.3
mms-1b-all 32.2/36.0 16.8/18.7 27.6/26.3 14.6/14.8 10.0/12.5 3.8/4.9 13.5/13.3 7.3/6.5 8.9/7.6 4.4/3.1
mms-1b-fl102 52.4/52.7 26.0/25.5 51.7/48.7 26.1/23.2 19.1/22.9 5.9/8.6 29.7/30.3 12.9/12.9 22.6/20.5 9.3/7.9
moonshine-base 26.4/26.2 18.1/17.5 17.0/13.8 11.6/9.1 6.4/7.5 3.4/3.9 5.8/7.0 3.4/4.1 11.7/9.9 6.7/4.6
moonshine-tiny 31.8/30.8 20.5/19.1 20.1/17.2 13.4/11.8 9.1/9.9 4.8/5.3 9.8/9.5 6.7/5.8 14.9/12.9 8.2/7.2
parakeet-ctc-0.6b 24.2/20.0 16.5/12.6 13.1/11.0 8.7/7.8 6.4/7.5 3.6/3.8 4.3/7.2 2.6/4.4 7.0/7.2 4.1/3.7
parakeet-ctc-1.1b 20.7/18.1 13.9/11.8 13.0/11.6 8.7/8.3 6.4/7.1 3.7/3.6 5.0/7.4 3.0/4.4 6.7/6.5 3.9/3.3
parakeet-rnnt-0.6b 21.9/17.9 15.4/12.2 14.5/11.6 10.0/8.6 4.9/7.3 2.9/3.7 5.0/6.9 3.0/4.0 6.4/6.7 3.8/3.7
parakeet-rnnt-1.1b 23.3/17.2 16.6/11.8 14.1/10.9 9.8/8.9 4.5/7.7 2.7/4.4 5.2/7.7 3.5/4.9 5.6/6.0 3.4/3.2
parakeet-tdt-1.1b 24.5/17.9 16.6/12.6 13.5/10.6 9.1/7.9 5.4/7.7 3.2/4.2 4.4/7.1 2.8/4.4 5.5/6.0 3.3/3.1
parakeet-tdt_ctc-110m 16.6/21.4 11.5/15.1 15.3/11.5 10.7/8.3 3.8/7.3 2.2/4.0 5.2/6.6 3.3/4.1 7.5/6.2 4.6/3.1
seamless-m4t-v2-large 35.0/36.3 25.1/24.9 45.1/43.2 35.9/34.7 11.7/13.6 7.6/8.7 26.5/25.5 21.0/19.1 8.0/7.8 5.7/5.0
speechllm-1.5B 45.1/44.5 32.7/32.0 60.3/60.8 44.1/46.8 10.6/10.9 6.2/5.8 19.4/17.6 14.2/11.9 30.8/29.9 22.0/21.3
speechllm-2B 52.9/53.1 36.6/35.7 36.9/37.8 25.8/27.3 14.6/15.3 8.1/7.5 18.8/16.7 12.9/9.4 28.7/27.7 18.5/17.5
stt_en_conformer_ctc_large 24.2/21.5 15.2/13.0 12.6/14.7 7.6/9.4 7.9/7.3 4.1/3.4 5.9/7.7 3.3/4.4 6.9/5.3 3.9/2.7
stt_en_conformer_ctc_small 31.3/26.9 19.0/15.7 16.6/17.8 9.6/11.7 10.0/9.4 5.1/4.1 8.0/9.9 3.9/5.2 8.9/7.3 5.0/3.8
stt_en_fastconformer_ctc_large 26.9/20.5 18.3/12.9 15.4/14.0 9.9/9.7 6.9/8.4 3.7/4.1 5.7/7.8 3.1/4.5 6.3/6.0 3.8/3.3
stt_en_fastconformer_transducer_large 26.5/20.4 19.0/13.8 16.9/15.1 11.3/11.1 6.0/7.9 3.4/4.0 4.9/7.0 2.8/4.4 6.7/6.9 4.1/3.8
wav2vec2-base-960h 44.7/40.1 24.5/20.6 27.3/28.7 13.5/15.7 21.5/22.4 8.9/8.7 13.8/14.7 6.1/6.6 20.5/19.4 9.1/7.8
wav2vec2-conformer-rel-pos-large-960h-ft 37.3/34.7 21.2/19.7 20.3/22.1 10.6/13.4 12.0/12.2 5.2/4.6 11.7/12.3 6.6/6.3 14.8/13.1 6.9/5.5
wav2vec2-conformer-rope-large-960h-ft 35.3/32.1 20.4/19.3 20.6/22.1 10.4/12.9 11.7/12.5 5.1/4.8 10.9/11.8 5.5/6.0 14.5/13.4 6.9/5.4
wav2vec2-large-960h 38.9/35.7 21.6/19.2 23.2/25.1 11.5/13.7 16.3/17.1 6.9/6.6 12.2/13.2 5.6/6.1 18.1/16.7 8.2/7.0
wav2vec2-large-960h-lv60-self 32.5/29.4 18.7/15.7 20.2/20.8 10.7/12.9 10.4/12.2 4.3/4.7 9.5/10.5 4.2/4.9 13.5/12.5 6.4/5.2
wav2vec2-large-robust-ft-libri-960h 36.2/32.6 19.2/16.9 20.9/23.0 9.6/12.5 11.8/12.5 5.0/4.8 10.6/11.9 4.8/5.4 15.4/14.0 7.0/5.8
whisper-large 31.2/32.4 24.6/23.0 17.6/12.6 13.7/10.5 3.7/7.4 2.1/4.4 19.3/16.1 14.0/10.3 8.9/6.0 5.5/3.2
whisper-large-v2 18.8/25.2 14.2/18.1 18.7/15.1 14.8/12.1 4.1/7.7 2.4/4.8 28.3/25.3 19.4/14.4 8.7/6.8 5.5/3.8
whisper-large-v3 20.4/27.5 15.6/19.5 15.6/11.9 11.8/8.7 3.2/6.3 1.7/4.0 10.5/8.3 8.7/5.5 11.0/8.6 7.8/6.1
whisper-large-v3-turbo 16.0/23.7 12.0/16.5 15.3/11.9 11.5/9.0 3.1/6.3 1.7/3.9 9.9/8.1 8.5/5.2 13.3/11.1 9.8/7.9
whisper-medium.en 20.1/25.1 15.3/18.0 21.2/16.1 16.6/13.2 4.0/7.7 2.2/5.0 17.3/14.6 18.3/14.2 9.0/7.2 5.6/3.2
whisper-small.en 21.2/25.1 16.5/18.0 18.2/14.1 13.9/10.9 3.9/7.5 2.1/4.6 10.6/8.3 15.6/12.0 9.5/8.1 5.9/3.4
whisper-tiny 30.1/31.9 21.7/21.5 24.0/20.3 17.5/14.4 8.1/11.9 3.9/6.7 17.6/15.3 13.0/9.5 13.2/11.2 7.4/6.3
Table 9: Actual and approximated WER and CER, separated by a forward slash, across five standard datasets. The regression model is trained on nine datasets and tested on one, with this process repeated for all datasets, ensuring that the test data is always out-of-distribution.
This is an appendix.